OpenStudy (anonymous):

(x5 + 7x3 – 9) + (4x5 – 2x3 + 13)

OpenStudy (anonymous):

add like terms ... x^5 + 4x^5 = 5x^5 Can you do the next two?

OpenStudy (angela210793):

add and subtract just as u do with real no.s one x^5+ four x^5=...... seven x^3- two x^3=...... -9+13=......